2 dice are rolled what are the odds that one will show an odd number and one will show an even number?
Answer by richard1234(7193)

We can either roll an odd followed by an even, or an even followed by an odd. Rolling an odd or an even occurs with probability 1/2, so

P(one odd and one even) = (1/2)(1/2) + (1/2)(1/2) = 1/2.

Or, we can find the probability of getting two odds or two evens, and subtract the result from 1.